OpenStudy (anonymous):

Which expression is equal to 6(b+4)

OpenStudy (owlcoffee):

Apply distributive to make it an equality: \[6(b+4)\] Then by the property: \[a(b+c)=ab+ac\] applying it: \[6(b+4)=6b+24\]